OpenStudy (anonymous):

PLZ PLZ HELP!! Medal given!! Write y = -2/5x - 3 in standard form using intergers

OpenStudy (anonymous):

a. 2x + 5y = -15 b. 2x + y = -3 c. 2x + 5y = -3 d. x + y = -15

OpenStudy (anonymous):

@rachelmcclellan @Papaluty

OpenStudy (anonymous):

d i think.

OpenStudy (anonymous):

Okay... Thanks